OpenStudy (jackierobinson12):

Solve the triangle. A = 19°, C = 102°, c = 6

OpenStudy (jackierobinson12):

@mathstudent55

OpenStudy (dumbcow):

First find angle B ..... sum of triangle is 180

OpenStudy (dumbcow):

Then use Law of Sines twice to find sides a and b \[\frac{a}{\sin A} = \frac{c}{\sin C}\]

OpenStudy (dumbcow):

\[\frac{b}{\sin B} = \frac{c}{\sin C}\]